At time t(in seconds), a particle of mass 3kg has position vector r (meter) where r=3ti^-4costj^.The impulse of the force during the time interval 0-t-900 is
Ans-12j^ NS

Grade:12th pass

1 Answers

nandhitha
4 years ago
m=3kg
r=3ti^-4costj^
v=dr/dt=ti^=sintj^
a=dv/dt=costj^
since,f=ma
f=(3kg)(costj^)
inpulse=delta p=(F)(delta t)
                        = (f)(t final – t initial)
                        =(3)(costj^)(t final – t initial)
you can simplify the rest